Parrott Corporation produced 4,000 units of their product in the current period. All work must be shown - Total 24 points Each unit is budgeted (Standard Cost) as follows: Direct materials, 5 yards at $ 3.50 per yard Direct labor, 7 hours at $ 8.00 per hour Factory overhead, $ 27 per unit produced. Actual costs for the current period’s production for 4,000 units are as follows: Direct materials, 21,500 yards at $ 3.30 per yard Direct labor, 27,900 hours at $8.10 per hour Factory overhead, $ 107,500 Directions: Calculate each of the following variances and indicate if favorable or unfavorable. All work must be shown - Total 24 points Direct materials variance – overall Direct materials quantity variance Direct materials price variance Direct labor variance – overall Direct labor time variance Direct labor rate variance Factory overhead variance Total variance – overall
Answer #1

Solution:

Direct Material Cost Variance
Actual Cost Standard cost for actual quantity Standard Cost
AQ * AP = AQ * SP = SQ * SP =
21500 $3.30 $70,950.00 21500 $3.50 $75,250.00 20000 $3.50 $70,000.00
$4,300.00 Favorable $5,250.00 Unfavorable
Direct Material Price Variance Direct Material Qty variance
Direct material price variance $4,300.00 Favorable
Direct material quantity variance $5,250.00 Unfavorable
Total direct material variance $950.00 Unfavorable
Direct Labor Cost Variance
Actual Cost Standard cost for actual quantity Standard Cost
AH * AR = AH * SR = SH * SR =
27900 $8.10 $225,990.00 27900 $8.00 $223,200.00 28000 $8.00 $224,000.00
$2,790.00 Unfavorable $800.00 Favorable
Direct Labor rate Variance Direct Labor Efficiency Variance
Direct Labor Rate variance $2,790.00 Unfavorable
Direct Labor Efficiency variance $800.00 Favorable
Total direct labor variance $1,990.00 Unfavorable

Factory overhead variance overall = Fixed overhead applied - Actual overhead = (4000*$27) - $107,500 = $500 Favorable